# LAuReL: Learned Augmented Residual Layer

### v3

@@ Abstract
One of the core pillars of efficient deep learning methods is architectural improvements, such as residual/skip connections, which have led to significantly better model convergence and quality. Since their introduction, residual connections have become ubiquitous not only in convolutional neural networks but also in transformer-based architectures, the backbone of LLMs.
In this paper, we introduce the Learned Augmented Residual Layer (LAuReL)—a novel generalization of the canonical residual connection—designed to serve as an in-situ replacement while outperforming it in both model quality and footprint metrics. Our experiments show that LAuReL can enhance model quality for both vision and language models while adding fewer parameters and incurring less latency and memory overhead than naively increasing parameter count.
@@ 1 Introduction
Model efficiency is of critical importance in the age of extremely large language and vision models. Even if a given model has impressive quality, its footprint metrics such as train-time compute required, inference latency, resident memory size, peak memory consumption, etc. dictate if it can be experimented with and/or deployed in real-world settings. A large and/or slow model, regardless of its performance on eval tasks, might be infeasible to train and serve, and might not be fit for applications where the user expects responsiveness.
LLMs such as Gemini 1.5 Flash (Gemini-Team et al., 2024), DeepSeek V3 (DeepSeek-A.I. et al., 2024) have been explicitly designed with these efficiencies in mind and consistently outperform much larger models that preceded them.
Consequently, improving the Pareto-frontier of model quality vs footprint, via efficient deep learning methods has been an area of active research in the past few years. Areas of interests span from algorithmic techniques (Menghani, 2023), to efficient hardware (Sze et al., 2017), to identifying best practices around model efficiency (Dehghani et al., 2022), etc.
One of the core pillars of efficient deep learning methods are architectural improvements such as the residual/skip connection, which had led to significantly better model convergence and quality (He et al., ). Since its introduction, the residual connection has become ubiquitous in not just convolutional neural networks but also transformer-based architectures (Vaswani et al., 2017), the backbone of LLMs of today.
@@
Here, $f(\cdot)$ can be any non-linear function such as attention, MLP, multiple non-linear layers, etc., $x_{i}$ is the input to the said non-linear function, and $x_{i+1}$ is the combined output of the non-linear function and the residual component. Refer to Figure 1 for an illustration. To simplify exposition, we ignore pre-processing functions such as layer norm, which can be folded into $f(\cdot)$ without loss of generality.
@@ 2 Learned Augmented Residual Layer
Here $\alpha$ is a learned scalar parameter, and $g(\cdot)$ is a learned linear function with $x_{i},x_{i-1},\ldots,x_{0}$ as inputs, where $x_{j}$ is the output of the $j$th residual connection. The intuition behind LAuReL is that one can learn a richer set of (linear) functions than just using $x_{i}$ as the residual component. One motivation behind seeking these richer linear functions is the concept of a “residual stream” (Elhage et al., 2021), where the residual connection is considered to be part of a stream of information that passes through each layer without being exposed to any non-linearities. This allows the learning process to focus on the non-linear components better.
Each layer / operation can read from, and subsequently write to this residual stream based on what it read. Given that the residual connection has been shown to be important for model quality and convergence, we designed LAuReL to operate on this residual stream in a learned fashion, while being light-weight in terms of the model size and latency changes.
@@ 2.1 Residual Weights Version (LAuReL-RW)
Notice that this version assigns learnable weights to the $f(x_{i})$ and $x_{i}$ from (1). In practice, we found that we cannot let $\alpha$ and $\beta$ grow unbounded, and using a normalization function such as softmax or sigmoid helps. Clearly, this version will add only two new parameters per LAuReL layer. If necessary, we can always replace these two parameters by a single learnable parameter and use a function like sigmoid to define $\alpha,\beta$ in terms of this single parameter to create a convex combination.
@@ 2.2 Low-Rank Version (LAuReL-LR)
In this version, we fix $\alpha=1$, and $g(x_{i})=Wx_{i}$ in (2) to obtain:
@@
Note that, as written, $W$ is a learnable $D\times D$ matrix, where $D$ is the model dimension for transformer-based models, or more generally it is the last dimension of $x_{i}$; hence the $W$ matrix will add $D^{2}$ new parameters (per LAuReL layer) to the model.
@@
Here, both $A$ and $B$ matrices are learnable. The number of new parameters is $2rD$.
@@ 2.3 Previous Activations Version (LAuReL-PA)
This version is similar to LAuReL-LR, except that we use use $k$ activations from the previous blocks. In particular, we set
$g(x_{i})=(\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}.$
@@
$x_{i+1}=f(x_{i})+(\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}.$ (4)
In practice, we replace $h_{i}$ by a low-rank product similar to the LAuReL-LR version, but using the identity function is also an option. When using the low-rank product for $h_{i}$, the number of new parameters is $2rD+k$, where $k$ is the number of previous activations we are accessing.
@@ 2.4 Other Derived Variants
Note that, all three proposed versions mentioned above are a combination of scalar and/or low-rank products on top of the vanilla residual connection in (1). This makes LAuReL especially light-weight in terms of its impact on model size and latency. We discuss the efficiency of LAuReL variants in more detail in Section 4.
That being said, our framework is generic enough to allow combinations of the above variants, as well as new variants. For instance, one straight-forward combination is LAuReL-RW+LR, where the residual weights from LAuReL-RW can be applied along with the LAuReL-LR variant to rewrite (3) as follows:
@@
$x_{i+1}=\alpha f(x_{i})+\beta((\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}))+x_{i}).$
@@
$x_{i+1}=\alpha f(x_{i})+\beta((\sum_{j=0}^{k-1}\gamma_{i,j}\cdot A_{i,j}B_{i,j}(x_{i-j}))+x_{i}).$ (6)
Yet another variant while slightly relaxing our above formulations would be to treat $\alpha$ and $\beta$ as vectors of length $D$, which is the size of the model dimension in case of transformer-like models. In other cases it is the size of the last dimension of the input $x_{i}$. In this variant we would learn fine-grained per-dimension weights when mixing $x_{i}$ and $f(x_{i})$ at the cost of $2D$ parameters per LAuReL-LR layer, instead of two per parameters per layer.
@@ 3 Experiments
We experiment with LAuReL in two domains, namely, vision and language. For the first case, our goal is to improve the image classification accuracy of the ResNet-50 model on the ImageNet-1K dataset (Deng et al., 2009). For the second case our goal is to improve the performance of two different large language models (LLMs) of size 1B and 4B params respectively, evaluated after the pre-training stage, on common benchmarks.
@@ 3.1 ResNet-50 on ImageNet-1K
In addition, we also find that if we simply add another layer to the ResNet-50 model (i.e., naive scaling), we can increase the model’s accuracy by $0.25\%$ to reach $75.20\%$, while adding $4.37\%$ new parameters. With that in context, applying LAuReL on the model leads to better results (see Table 1).
If we only use the LAuReL-RW version, we get an improvement of $0.15\%$ on average with only $0.003\%$ extra parameters, which is essentially negligible. When we try the LAuReL-RW+LR version from (5) with $r=16$, we achieve an accuracy of $75.20\%$ while adding only $1.68\%$ extra parameters; this matches the performance of the baseline with an extra layer, while using 2.6×2.6$\times$2.6 × fewer extra parameters. Additionally, when we use the combined LAuReL-RW+LR+PA version from (6) we improve the accuracy to $75.25\%$ while still using 1.82×1.82$\times$1.82 × fewer extra parameters than the baseline with one extra layer, demonstrating that LAuReL is superior to naively scaling the model. Notably even though we make fundamental changes to the residual connection we did not find any training instabilities when using LAuReL.
Table 1: Applying LAuReL on a ResNet-50 trained on the ImageNet 1K classification dataset. LAuReL-RW provides a significant boost with negligible extra parameters. LAuReL-RW+LR and LAuReL-RW+LR+PA meet and beat the naive scaling baseline while using 2.6×2.6$\times$2.6 × and 1.82×1.82$\times$1.82 × fewer parameters.
@@ 3.2 Large-Scale LLM Pre-training
In this setup, our goal was to test the performance of LAuReL when applied on top of strong LLMs. During the course of our work we evaluated LAuReL on two separate LLMs, which we pre-trained from scratch. The first LLM (LLM-1) was a 1B parameter model pre-trained with a data-mixture consisting of text tokens only. The second LLM (LLM-2) was a 4B parameter model which was pre-trained with a multi-modal and multi-lingual data mixture. Both LLMs were trained with $\Theta(0.5T)$ tokens.
What we varied across the two LLMs was to allow for different budgets for increasing model footprint (params, latency, memory, etc.) when applying LAuReL. For LLM-1 we allow a very small increase in these metrics ($\sim$ 0.01% extra params, and nearly no latency increase). For LLM-2 we allow a slightly lenient, yet very modest increase in params ($\sim$ 0.1% extra params), and latency (1% increase).
Given the scale of LLMs today, both the budgets would be considered negligible. For instance, a $0.1\%$ increase in parameters for a 4B model corresponds to 4M more parameters. As demonstrated in the ResNet / ImageNet experiments in Section 3.1, LAuReL outperforms naive scaling, but we further detail the comparison with naive scaling in Section 4.4.
Our objective behind testing LAuReL in these conditions was to demonstrate the efficacy of the method and ensure that LAuReL scales well across different LLM setups in the wild.
3.2.1 LLM-1: 1B params, $\Theta(0.5\textrm{T})$ tokens, very low extra params/latency/memory budget
For our first baseline, we chose a 1B parameter decoder-only model based on the Transformer architecture. We pre-trained both the baseline, and our experiment with LAuReL, from scratch; we use the LAuReL-RW and LAuReL-LR versions (with $r=4$). Both the models were trained using 256 Google Cloud TPU v5e chips for approximately two weeks each, using a pre-training mixture consisting of only text data which included webpages, books, code, translations, and other text data.
It is worth noting that the combined LAuReL-RW+LR variant adds only 0.012% more parameters as compared to the baseline model. Since we chose $r=4$, the total number of parameters added by LAuReL-LR is $8ND$ and $2N$ parameters by LAuReL-RW, for a total of $2N(4D+1)$ parameters. Typically $N\in[10,100]$ and $D\in[500,5000]$. For the sake of illustration, assume $N=20$ and $D=1000$, using the LAuReL-RW+LR leads to $160,040$ extra params in a 1B param model. Thus, the number of new parameters is dwarfed by that of the original model. Additionally, we found that the additional latency introduced by LAuReL-RW+LR was within the noise range.
@@
Table 2: Evaluation results of LLM-1 as described in Section 3.2.1; a 1B parameter Decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common eval benchmarks (higher is better for all task type and task combinations listed below). The LAuReL variant outperforms the baseline on all but one dataset while adding only 0.012% extra parameters.
@@
Table 3: Evaluation results of LLM-2 as described in Section 3.2.2; a 4B parameter Decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common eval benchmarks (higher is better for all task type and task combinations listed below). The LAuReL variant outperforms the baseline on all but two tasks, while adding only $\sim$ 0.1% extra parameters.
@@
3.2.2 LLM-2: 4B params, $\Theta(0.5\textrm{T})$ tokens, low extra params/latency/memory budget
In this second setting, we worked on a bigger 4B parameter decoder-only model with a similar token budget, but trained on a multimodal and multilingual corpus of tokens.
To compensate for a 4×4$\times$4 × larger model, and also the fact that the dataset and eval tasks are harder, we allowing a bigger footprint budget for LAuReL; $\sim 0.1\%$ extra parameters and $\sim 1\%$ extra latency. As mentioned earlier, this is still a negligible increase in model params and latency.
@@
In this setting, both the baseline and the LAuReL experiment were trained using 1024 Google Cloud TPU v4 chips for slightly more than two days each. See Table 3 for the evaluation results of both the baseline model and the model with LAuReL. The task type and individual tasks are listed in the first and second columns respectively, and a higher score is better for all the tasks. LLM-2 had a sophisticated suite of evaluation tasks, including math, general reasoning, reading comprehension, translation, and multimodal tasks. All of the listed eval tasks are used by leading LLMs to evaluate model quality, further solidifying our confidence in the model evaluation setup.
To start off, two of the LLM-2 evaluation tasks, Math, and MMLU were common with LLM-1 (refer to Table 2 for LLM-1 results). It can be seen that LLM-2 is much stronger than LLM-1 on both the common evals; 4.16×4.16$\times$4.16 × better on the Math task, and 1.93×1.93$\times$1.93 × better on the MMLU task. We attribute this to LLM-2 being 4×4$\times$4 × larger, and having a more sophisticated pre-training mixture. LAuReL using the LAuReL-RW+LR variant improves on both Math and MMLU, which is pleasantly surprising given how powerful is LLM-2. For other tasks LAuReL improves significantly over the baseline as well, except for WMT23 and DocVQA, where it was neutral.
In terms of costs, the model adds $\sim 0.1\%$ more parameters. This is still a very reasonable approximately 4M parameters on top of a 4B parameter model. In terms of latency, we tested on both server-side (Google CloudTPU) for cloud serving, and a leading smartphone for on-device inference. In both the benchmarks, we measure nearly 1-2% increase in latency for prefill and generation. There was no human perceptible difference in terms of time-to-first-token.
We did not try the LAuReL-PA version for the above LLM experiments, as the LLM training was expensive. Additionally the LAuReL-PA variant requires slightly more memory to keep previous block activations accessible. This also meant changing sharding, rematerialization, and other settings in the LLM codebases we worked on. However, we expect the LAuReL-PA results from the ResNet experiments to also hold in this case as well.

### v4

@@ Abstract
One of the core pillars of efficient deep learning methods are architectural improvements, such as residual/skip connections, which have led to significantly better model convergence and quality. Since their introduction, residual connections have become ubiquitous not only in convolutional neural networks but also in transformer-based architectures, the backbone of LLMs.
In this paper, we introduce the Learned Augmented Residual Layer (LAuReL)—a novel generalization of the canonical residual connection—designed to serve as an in-situ replacement while outperforming it in both model quality and footprint metrics. Our experiments show that LAuReL can enhance quality for both vision and language models while adding fewer parameters and incurring less latency and memory overhead than naively increasing parameter count.
@@ 1 Introduction
Model efficiency is of critical importance in the age of extremely large language and vision models. Even if a given model has impressive quality, its footprint metrics such as train-time compute, inference latency, resident memory size, peak memory consumption, etc. dictate if it can be experimented with and/or deployed in real-world settings. A large and slow model can be impractical to train and use, making it unsuitable for applications that require fast responses, no matter how well it performs on benchmarks.
LLMs such as Gemini 1.5 Flash (Gemini-Team et al., 2024), DeepSeek V3 (DeepSeek-A.I. et al., 2024) have been explicitly designed with these efficiencies in mind and consistently outperform much larger models that preceded them. Consequently, improving the Pareto-frontier of model quality and model footprint, via efficient learning methods has been an area of active research in the past few years. Areas of interests span from algorithmic techniques (Menghani, 2023) and efficient hardware (Sze et al., 2017) to best practices around model efficiency (Dehghani et al., 2022).
One of the core pillars of efficient deep learning methods are architectural improvements such as the residual/skip connection, which had led to significantly better model convergence and quality (He et al., ). The residual connection has become ubiquitous not only in convolutional neural networks but also in transformer-based architectures (Vaswani et al., 2017), which are the backbone of today’s LLMs.
@@
Here, $f(\cdot)$ can be any non-linear function such as attention, MLP, multiple non-linear layers, etc., $x_{i}$ is the input to the said non-linear function, and $x_{i+1}$ is the combined output of the non-linear function and the residual component (Figure 1). For simplicity, we ignore pre-processing functions such as layer norm, which can be folded into $f(\cdot)$.
@@ 2 Learned Augmented Residual Layer
Here, $\alpha$ is a learned scalar parameter, and $g(\cdot)$ is a learned linear function with $x_{i},x_{i-1},\ldots,x_{0}$ as inputs, where $x_{j}$ is the output of the $j$-th residual connection.
The main intuition is that one can learn a richer set of (linear) functions than just using $x_{i}$ as the residual component. One motivation behind seeking these richer linear functions is the concept of a “residual stream” (Elhage et al., 2021), where the residual connection is considered to be part of a stream of information that passes through each layer without being exposed to any non-linearities. This allows the learning process to focus on the non-linear components better.
Each layer/operation can read from, and subsequently write to this residual stream. Since the residual connection has been shown to be important for model quality and convergence, we designed LAuReL to operate on this residual stream in a learned fashion, while being light-weight in terms of the model size and latency changes.
@@ 2.1 Residual Weights Version (LAuReL-RW)
Notice that this version assigns learnable weights to the $f(x_{i})$ and $x_{i}$ from (1). In practice, we found that we cannot let $\alpha$ and $\beta$ grow unbounded, and using a normalization function such as softmax or sigmoid helps. Clearly, this version will add only two new parameters per LAuReL layer. If necessary, we can replace these two parameters by a single learnable parameter and use a function such as sigmoid to define $\alpha,\beta$ in terms of this single parameter.
This variant can be useful for learning the relative importance of the non-linear component ($f(x_{i})$) and the residual input ($x_{i}$). In the earlier layers the former might be more important, while in the later layers, the latter could be useful for mitigating problems such as vanishing gradients. This variant can help by adaptively learning these weights.
@@ 2.2 Low-Rank Version (LAuReL-LR)
Here, we fix $\alpha=1$, and $g(x_{i})=Wx_{i}$ in (2) to obtain:
@@
As written, $W$ is a learnable $D\times D$ matrix, where $D$ is the model dimension for transformer-based models, or more generally it is the last dimension of $x_{i}$. Hence, the $W$ matrix will add $D^{2}$ new parameters (per LAuReL layer).
@@
Here, both $A$ and $B$ matrices are learnable. The number of new parameters is $2rD$, per LAuReL layer.
This variant helps with allocating learning capacity for the linear part of the network (the residual input, i.e., $x_{i}$ in Figure 2), such that the main network can use its capacity towards learning better nonlinear functions ($f(x_{i})$), while LAuReL contributes the linear components ($x_{i}+BAx_{i}$) to the residual stream.
@@ 2.3 Previous Activations Version (LAuReL-PA)
This is similar to LAuReL-LR, except that we use $k$ activations from the previous blocks. In particular, we set
$g(x_{i},\ldots,x_{0})=x_{i}+\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}),$
@@
$x_{i+1}=f(x_{i})+x_{i}+\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j}).$ (4)
In practice, we replace $h_{i}$ by a low-rank product similar to the LAuReL-LR version, but using the identity function is also an option. When using a rank $r$ product for $h_{i}$, the number of new parameters per LAuReL layer is $2rD+k$, where $k$ is the number of previous activations used.
We can consider this variant to be a hybrid of the LAuReL-RW and LAuReL-LR variants, where multiple previous activations are used in a weighted manner that is learned. This allows layers accelerated access to previous activations, along with learning their relative importance.
@@ 2.4 Other Derived Variants
All three proposed LAuReL versions are a combination of scalar and/or low-rank products on top of the vanilla residual connection in (1). This makes it especially light-weight in terms of its impact on model size and latency. We discuss the efficiency of the variants in more detail in Section 4.
That being said, LAuReL is generic enough to allow combinations of the above variants, as well as new variants. For instance, one straight-forward combination is LAuReL-RW+LR, where the residual weights from LAuReL-RW can be applied along with the LAuReL-LR variant to rewrite (3) as follows:
@@
$x_{i+1}=\alpha f(x_{i})+\beta\left(x_{i}+\sum_{j=0}^{k-1}\gamma_{i,j}\cdot h_{i}(x_{i-j})\right).$
@@
$x_{i+1}=\alpha f(x_{i})+\beta\left(x_{i}+\sum_{j=0}^{k-1}\gamma_{i,j}\cdot A_{i,j}B_{i,j}(x_{i-j})\right).$ (6)
Yet another variant while slightly relaxing our above formulations would be to treat $\alpha$ and $\beta$ as vectors of length $D$. Here, we would learn fine-grained per-dimension weights when mixing $x_{i}$ and $f(x_{i})$ at the cost of $2D$ parameters per LAuReL-LR layer, instead of two per parameters per layer.
@@ 3 Experiments
We experiment with LAuReL in two domains, namely, vision and language. For the first case, our goal is to improve the image classification accuracy of the ResNet-50 model on the ImageNet-1K dataset (Deng et al., 2009). For the second case our goal is to improve the performance of two different large language models (LLMs) of size 1B and 4B parameters respectively, evaluated after the pre-training stage, on common benchmarks.
@@ 3.1 ResNet-50 on ImageNet-1K
In addition, we also find that if we simply add another layer to the ResNet-50 model (i.e., naive scaling), we can increase the model’s accuracy by $0.25\%$ to reach $75.20\%$, while adding $4.37\%$ new parameters. With that in context, applying LAuReL on the model improves it (Table 1).
If we only use the LAuReL-RW version, we get an improvement of $0.15\%$ on average with only $0.003\%$ extra parameters, which is essentially negligible. When we try the LAuReL-RW+LR version from (5) with $r=16$, we achieve $75.20\%$ accuracy while adding only $1.68\%$ new parameters; this matches the performance of the baseline with an extra layer, while using 2.6×2.6$\times$2.6 × fewer extra parameters. Additionally, when we use the combined LAuReL-RW+LR+PA version from (6), we improve the accuracy to $75.25\%$ while still using 1.82×1.82$\times$1.82 × fewer extra parameters than the baseline with one extra layer, demonstrating that LAuReL is superior to naively scaling the model. Notably, despite substantial changes to the residual connection, we did not find any training instabilities with LAuReL.
Table 1: Applying LAuReL on a ResNet-50 trained on the ImageNet 1K classification dataset. LAuReL-RW provides a significant boost with negligible extra parameters. LAuReL-RW+LR and LAuReL-RW+LR+PA meet and beat the naive scaling baseline while using 2.6×2.6$\times$2.6 × and 1.82×1.82$\times$1.82 × fewer parameters. Results that provide statistically significant boost over the baseline are highlighted in green and bold.
@@ 3.2 Large-Scale LLM Pre-training
In this setup, our goal was to test the performance of LAuReL when applied on top of strong LLMs. During the course of our work we evaluated LAuReL on two separate LLMs, which we pre-trained from scratch. The first LLM (LLM-1B) is a 1B parameter model pre-trained with a data-mixture consisting of only text tokenss. The second LLM (LLM-4B) is a 4B parameter model that was pre-trained with a multi-modal and multi-lingual data mixture. Both LLMs were trained with $\sim 0.5$T tokens.
What we varied across the two LLMs was to allow for different budgets for increasing model footprint (parameters, latency, memory, etc.) when applying LAuReL. For LLM-1B we allow a very small increase in these metrics ($\sim$ 0.01% extra parameters, and nearly no latency increase). For LLM-4B we allow a lenient, yet modest increase in parameters ($\sim$ 0.1% extra parameters), and latency (1% increase).
Given the scale of LLMs today, both the budgets would be considered negligible. For instance, a $0.1\%$ increase in parameters for a 4B model will only correspond to 4M more parameters. As demonstrated in the ResNet-50/ImageNet experiments (Section 3.1), LAuReL outperforms naive scaling; see Section 4.5 for a more detailed comparison.
Our objective behind testing LAuReL in these conditions was to demonstrate its efficacy and ensure that it scales well across different LLM setups in the wild.
3.2.1 LLM-1B: Very low additional footprint
For our first baseline, we chose a 1B parameter decoder-only transformer-based model. We pre-trained both the baseline, and our experiment with LAuReL, from scratch; we use the LAuReL-RW and LAuReL-LR versions (with $r=4$). Both the models were trained using 256 Google Cloud TPU v5e chips for approximately two weeks each, using a pre-training mixture consisting of only text data that included webpages, books, code, and translations.
It is worth noting that the combined LAuReL-RW+LR variant adds only 0.012% more parameters as compared to the baseline model. Since we chose $r=4$, the number of parameters added by LAuReL-LR is $8ND$ and the number of parameters added by LAuReL-RW is $2N$, for a total of $2N(4D+1)$ additional parameters. Typically $N\in[10,100]$ and $D\in[500,5000]$. For the sake of illustration, assuming $N=20$ and $D=1000$ and using LAuReL-RW+LR leads to $160,040$ extra parameters in a 1B parameter model. Thus, the number of new parameters is dwarfed by that of the original model. Furthermore, the additional latency introduced by LAuReL-RW+LR was within the noise range.
@@ 3.2.1 LLM-1B: Very low additional footprint
Table 2: Evaluation results on LLM-1B as described in Section 3.2.1; a 1B parameter decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common evaluation benchmarks (higher is better for all task type and task combinations listed below). LAuReL variant outperforms the baseline on all but one dataset while adding only 0.012% extra parameters. Results that provide $\geq 2\%$ relative improvement over the baseline are highlighted in green and bold.
@@
Table 3: Evaluation results on LLM-4B as described in Section 3.2.2; a 4B parameter decoder-only LLM pre-trained from scratch with (a) once with the baseline model architecture, and (b) once using LAuReL on top. We evaluated both the models on a number of common evaluation benchmarks (higher is better for all task type and task combinations listed below). LAuReL variant outperforms the baseline on all but two tasks, while adding only $\sim$ 0.1% extra parameters. Results that provide statistically significant improvement over the baseline are highlighted in green and bold.
@@
3.2.2 LLM-4B: Low additional footprint
In this second setting, we experimented with a 4B parameter decoder-only model with a similar token budget, but trained on a multimodal and multilingual corpus of tokens.
To compensate for a 4×4$\times$4 × larger model, and also the fact that the dataset and the evaluation tasks are harder, we allowing a bigger footprint budget for LAuReL; $\sim 0.1\%$ extra parameters and $\sim 1\%$ extra latency. Note that this is still a negligible increase in model parameters and latency.
@@ 3.2.2 LLM-4B: Low additional footprint
In this setting, both the baseline and the LAuReL experiment were trained using 1024 Google Cloud TPU v4 chips for slightly more than two days each. See Table 3 for the evaluation results of both the baseline model and the model with LAuReL. The task type and individual tasks are listed in the first and second columns respectively, and a higher score is better for all the tasks. LLM-4B had a sophisticated suite of evaluation tasks, including math, general reasoning, reading comprehension, translation, and multimodal tasks. All of the listed evaluation tasks are used by leading LLMs to evaluate model quality, further solidifying our confidence in the model evaluation setup.
To start off, two of the LLM-4B evaluation tasks, Math, and MMLU were common with LLM-1B (refer to Table 2 for LLM-1B results). It can be seen that LLM-4B is much stronger than LLM-1B on both the common evals; 4.16×4.16$\times$4.16 × better on the Math task, and 1.93×1.93$\times$1.93 × better on the MMLU task. We attribute this to LLM-4B being 4×4$\times$4 × larger, and having a more sophisticated pre-training mixture. LAuReL using the LAuReL-RW+LR variant improves on both Math and MMLU, which is pleasantly surprising given how powerful is LLM-4B. For other tasks LAuReL improves significantly over the baseline as well, except for WMT23 and DocVQA, where it was neutral.
In terms of costs, the model adds $\sim 0.1\%$ more parameters. This is still a very reasonable since it means additional 4M parameters on top of a 4B parameter model. In terms of latency, we tested on both server-side (Google CloudTPU) for cloud serving, and a leading smartphone for on-device inference. In both the benchmarks, we measure nearly 1–2% increase in latency for prefill and generation. There was no human perceptible difference in terms of time-to-first-token.
We did not try the LAuReL-PA version for the above LLM experiments, as the LLM training was expensive. However, we expect the LAuReL-PA results from the ResNet experiments to also hold in this case as well.